Status Report

ISS On-orbit Status 30 Aug 2002

By SpaceRef Editor
August 30, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

The crew had a normal workday.  Tomorrow: begin of their on-orbit Labor Day weekend.

CDR Valeriy Korzun and FE-2 Sergey Treschev completed final post-EVA close-out operations.  [These included removal and stowage of the portable BNP O2 (oxygen) repressurization tanks from the Service Module (SM) Transfer Compartment (PkhO) and DC-1 Docking Compartment, returning the DC-1 to a nominal configuration, and stowing bagged tools and hardware, such as hammer, screwdriver, wire cutters, wrenches, tool carrier (KPU), transport tethers, etc.  Stowage locations are in the DC-1, SM, US Airlock (A/L), FGB and behind a Node panel.]

The cosmonauts also started gathering the necessary inventory of tools, cables and cable adapters for next week’s major reconfiguration of the Russian onboard Ethernet crew support network BVS (similar to the U.S. OpsLAN).  [This will involve dismantling the existing SM coaxial network and connecting the SSC3 (station support computer #3) to the BRI SmartSwitch router which Mikhail Tyurin tested in November last year.  The actual routing of the cables was planned for 9/3 (Tuesday), followed by the full implementation with the SSC on 9/5 (Thursday).  MMT (Mission Management Team) has requested to delay routing and implementation until the reconfiguration has been fully concurred in by both sides, including station crew and 5S visiting crew.  Final resolution will be reviewed by MMT on 9/3.]

As part of the ongoing troubleshooting of the A/L BSA’s (battery storage assembly) charging/discharging capability, FE-1 Peggy Whitson initiated draining of two other EVA batteries by running the EMU (extravehicular mobility unit) backpack fan.  The discharged batteries will then be used to work with the BSA.

In concluding this week’s onboard water sampling, Peggy performed the analysis of the MCDs (microbial capture devices) after any contained microbial colonies had two days for growing in the samples taken on 8/28.  The results were entered into the MEC (medical equipment computer).

Later, she power-cycled the IVCPDS (intravehicular charged particle directional spectrometer) after calling down its current display readings.  When the unit was powered up again, she once more called down the display data.

FE-2 Treschev terminated the regeneration process for absorption bed #2 of the BMP micropurification unit.  Both filter channels are now back in Purify mode.

Later, Sergey worked about an hour in the SM checking out the function of a recently installed ventilation fan (VV2), behind a wall panel inside an air duct.

Daily routine maintenance was completed by Korzun (SOSh life support systems servicing), Whitson (status checkup of PCG-STES008 and ADVASC payloads), and Treschev (inventory management system delta file preparation).

There will be no MSG (Microgravity Science Glovebox) activity this weekend.  After the MSG clean-up by Peggy Whitson, the PSRP (Payload Safety Review Panel) was to discuss today the resumption of MSG/SUBSA operations (solidification using a baffle in sealed ampoule).  After approval is obtained, there is still some work to be done by the SUBSA team, so that the earliest date for getting a sample started in MSG will be 9/2 (Monday).  SUBSA will request additional science runs to be added to the currently remaining six runs prior to stowage on 9/30.

Removal and replacement of a smoke detector (SD) in Express Rack 2 (ER2) was added to the task list, at the crew’s disposition. [The SD data has been “out of family” for the last couple of weeks, and MCC-H wants to avoid any risk to 25 days of continuous ZCG (Zeolite crystal growth) furnace operations during the 9A stage. Since ZCG uses the rack air for cooling, it is essential that ER2 has a healthy SD to support these operations.  The replacement task involves a rack rotation, which takes two crewmembers 90 min each way, and the SD R&R itself takes a single crewmember 45 min.  Today the crew could not locate a tool required for the ER2 rotation.]

A second item added to the “job jar” task list is to restart the MCA (major constituent analyzer), which requires a pumpout of its vacuum chamber.  [The MCA in the AR (atmosphere revitalization) rack, delivered on 5A with the Lab, is a mass spectrometer that draws ISS cabin air into a vacuum, ionizes it with the filaments and then produces ion trajectories which ?give away? major constituents by their (known) trajectories.  The instrument, which is similar to the Russian gas analyzer (GA) in the SM, has not been working since last year.]

After the vacuum exposure test of the CDRA (carbon dioxide removal assembly) catalyst bed on 8/28 (to characterize its leak rate), start-up of the unit is scheduled for next week, since it has to wait until the CCAA (common cabin air assembly) air conditioner, currently on its starboard channel, has been switched to port.

PL1 MDM (payload computer #1) data dumps are being studied at present to determine the cause of its recent failure to Diagnostics mode.  PL1 is currently in good backup state, with PL2 in Primary function.

Preliminary finding after yesterday’s spin testing of the spare CMG (control moment gyroscope) is positive.  MCC-H specialists are confident that the unit can be flown on ULF-1/STS-114 as planned.

After the failure yesterday of subset #3, the SM TsVM (Central Computer system) continues to run nominally on the two remaining lanes. Moscow plans to let it continue in this state until a second channel fails, before reinitializing the triple-redundant set.

After approximately 100 kg of UDMH fuel was transferred from Progress 8P to SM fuel tank #1 without problem on 8/28, yesterday’s transfer of N2O4 oxidizer was interrupted by failure of the transfer pump.  TsUP identified the problem as a programming bug in the top-level software (i.e., not a technical or hardware issue), made the necessary change and today completed the propellant transfer successfully.

A longer-term issue being worked jointly by U.S. and Russian experts continues to be the development of a flight attitude timeline/schedule which is acceptable to both sides for the critical September/October period (Russia favoring XPOP, US favoring LVLH).  [For their planned testing of the SM Kurs-P rendezvous radio system on 9/23 (prior to Progress 9P arrival), TsUP will probably again require XPOP as the last time.  Flight attitude must also be balanced between providing sufficient core power for the US segment (USOS) and keeping the temperatures of the docked Progress below 35 degC.  One acceptable flight attitude found to date involves XVV (x-axis in velocity vector, i.e., facing forward), with a 90 deg roll angle towards the Sun.]

Also jointly in work is the issue of optimal crew sleep shifting for the upcoming arrivals of Progress 9P and Soyuz 5S.

TVIS engineers at MCC-H continue to troubleshoot the data transfer problem from the TVIS PCMCIA (Personal Computer Memory Card International Association) to the MEC (medical equipment computer) for downlink to Earth.  Russian crew health specialists are particularly concerned about getting the results of cosmonaut exercising on 9/4 and 9/5 down.  TsUP is also considering to fly a treadmill from the old Buran program on Progress 9P as a TVIS backstop.

The Flexpack canisters of the RED (resistive exercise device) will reach the end of their certified lifetime next week (9/4).  Work is underway to extend the certification.

Russian engineers continue to work on the issue of the air leak between SM PkhO and FGB during the EVA-8 depressurization phase.  No word yet on findings.

TsUP has scheduled a reboost of the ISS by Progress 8P on 9/13 (Friday).  Planned velocity change is 2 m/sec.

Results of the failure investigation (by CSA’s Canadarm2 contractor, MD-R in Brampton) of the SSRMS Wrist Roll Joint (WRJ) brought back on UF-2 were presented yesterday at JSC.  Cause of the short circuit in the WRJ was found to have been a tiny piece of wire (0.25 inch long, human hair thickness) in a connector, which had escaped inspection.

Launch date of Progress 8P continues to be 9/20.  With docking planned for 9/24, ascent and rendezvous phase will be two days longer than nominal, for the purpose of conducting a Kurs-A system test during the fly-alone period.

As part of the Russian health maintenance program, the two cosmonauts were advised to start taking preventative cardiotropic medication tomorrow (i.e., drugs intended to prevent possible heart muscle shrinkage).

Shuttle Launch Schedule:

STS-112/9A –10/2

STS-113/11A — 11/10

STS-107 — NET 1/16/03

Today’s targets of the CEO (crew earth observations) program were Taiwan Smog (crew to look left and right of track), Eastern Mediterranean Dust (plumes continue to exit from north Africa, now also moving over the Nile delta.  Scientists concluded recently that the raw material of soils in the Middle East is mainly Saharan dust!), Congo-Zimbabwe Biomass Burning (crew to look left and right of track for fires and smoke-plume trajectories), and Lower Amazon River Basin (crew was to look right for detailed views of the estuary of the largest river on the planet. Shoreline changes, especially of islands, are of special interest. The science of sedimentology does not allow for the scale of movements recently documented here).
The ISS crew’s comparative images of the East European floodings taken 8/18 and 8/23 (high and receding flood stages) of the Danube floodplain at Budapest are being prepared for NASA?s “Earth Observatory” website for release within a few days.
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:19 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32-amp mode), on backup pump.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber on command override.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 761, temperature (deg C) — 26.9, ppO2 (mmHg) — 146.5, ppCO2 (mmHg) — 2.7.
  • SM Transfer Compartment:  Pressure (mmHg) — 751, temperature (deg C) — 21.4.
  • FGB Cabin:  Pressure (mmHg) — 752, temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 752.16, temperature (deg C) — 25.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 753.6, temperature (deg C) — 24.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 753.6, temperature (deg C) — 26.0; shell heater temp (deg C) — 25.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.4
  • PMA-2:  Shell heater temp (deg C) — 13.7

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] was failed and is now in Idle mode, requiring pumpdown of its vacuum chamber). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B in AutoTrack mode (solar-tracking).
  • SM batteries:   Battery #1 is off-line; all other batteries (7) are in “Complete Charge” mode.
  • FGB batteries:  Battery #5 is off-line; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is Off; SKV-2 is On.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off (cold backup); PL-2 MDM is operating as primary.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line.

  • State vector — US GPS (SIGI string 1)
  • Attitude — Russian segment
  • Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-8 viewing position, with Keep Alive power on both strings (based on MBS).
  • MBS: Keep Alive power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 6:58am EDT [= epoch]):

  • Mean altitude — 391.0 km
  • Apogee — 401.9 km
  • Perigee — 380.0 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0016206
  • Orbits per 24-hr. day — 15.59
  • Solar Beta Angle — -35.6 deg (magnitude decreasing)
  • Altitude decrease — 120 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 21561
  • Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane = ‘sun-fixed’ [yaw: -0.5 deg, pitch: -5.7 deg., roll: 0 deg]). Will remain in XPOP until solar Beta drops below -10 deg on 9/6.

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.